OpenStudy (anonymous):

Enthalpy change of formation of cyclopropane : +32.2 Enthalpy change of atomisation of graphite : +717 C-H bond energy : 410 H-H energy : 436 Find the bond energy of C-C in cyclopropane

OpenStudy (abmon98):



OpenStudy (abmon98):

6*C-H

OpenStudy (abmon98):

6*410=2460

OpenStudy (abmon98):

enthalpy of bonds broken+enthalpy of bond released 2460+(1*C-C)+0=2118 C-C=-342

OpenStudy (anonymous):

Ah thanks, how did your paper 1 go?

OpenStudy (abmon98):

is my working right

OpenStudy (abmon98):

paper 1 is 3 more hours to go iam worried about it

OpenStudy (anonymous):

Which variant?

OpenStudy (abmon98):

variant 1